OpenStudy (anonymous):

A principal of $450 is deposited in an account that pays 2.5% interest compounded yearly. Find the account balance after 3 years rounded to the nearest cent.

OpenStudy (johnweldon1993):

\[\large A = P(1 + \frac{ r }{ n })^{nt}\] P = principal r = rate in decimal form n = number of times per year the interest is compounded t = time in years Can you solve this?

OpenStudy (anonymous):

is the r 2.5 ?

OpenStudy (johnweldon1993):

Remember that is the percent form...to get the decimal form...put the number over 100 \[\large \frac{2.5}{100} \] This comes out to.....\(\large .025 \)

OpenStudy (anonymous):

453.75 ??

OpenStudy (johnweldon1993):

\[\large A = P(1 + \frac{ r }{ n })^{nt}\] \[\large A = 450(1 + \frac{ .025 }{ 1 })^{1 \times 3}\] \[\large A = 450(1.025 )^{3}\] \[\large A = 450(1.076890625)\] \[\large A = 484.6\] Is what I get...see what I did?

OpenStudy (anonymous):

ohhh i get it now !!! is this rounded ?

OpenStudy (anonymous):

@johnweldon1993 ????

OpenStudy (johnweldon1993):

That is rounded to the nearest cent as your question asks yes

OpenStudy (anonymous):

thanks.. how would you do this Membership is a local tennis club can be represented by the model, where t is the number of years since 2009. What was the membership cost in 2011 rounded to the nearest cent?